OpenStudy (anonymous):

How do I find the radius of a sector formed on a circle of central angle 1/4 radians and area 6 square inches?

OpenStudy (campbell_st):

well the area of the sector formula is \[A = \frac{\theta}{2} \times r^2\] where theta is the sector angle.. so substituting you have \[6 = \frac{\frac{\pi}{4}}{2} \times r^2\] just solve for r

OpenStudy (anonymous):

Okay, this is what I have right now: 6=1/2(2pie/4)(r^2)

OpenStudy (campbell_st):

oops my mistake the angle says 1/4 radians.... i read it as pi/4 so 6 = 1/2 x 1/4 times r^2 so 6 = 1/8 r^2 or r^2 = 48 solve for r...

OpenStudy (anonymous):

Okay if I did it correctly r is 4 root 3

OpenStudy (campbell_st):

yes that would be correct